# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 5.3 Abnormal grain growth
After systematically benchmarks and knowledge gained from quasi-static single GB cases, we investigate different abnormal grain growth patterns of materials with different defect segregation energies $\Delta g_{\text{dop}}$, $\Delta g_{\text{V}_{\text{O}}^{..}}$ and bulk defect concentrations such as $c_{\text{dop},\text{b}}$. The segregation of point defects in the core plays a significant role in influencing grain growth kinetics. To investigate the grain size distribution of STO polycrystalline material after sintering, we conduct 2D phase-field simulations. Note that three-dimensional phase-field simulations of the proposed model are achievable but require significantly more computational resources, which will be considered in the near future. In fact, to reveal the fundamental mechanisms, 2D simulations should be sufficient and allow for a comprehensive parameter study given the limited computational resources.
@@
The simulation box size is set to 390 nm $\times$ 390 nm with a grid resolution of 1000 $\times$ 1000 points. The polycrystalline microstructure is initialized using a Voronoi tessellation method, where each grain was assigned a unique grain index. We started with an initial number of 1000 grains, and the random seed used for generating the polycrystalline microstructure is set to 42. To optimize memory usage during simulations, we also utilize grain tracker algorithm [34]. In the simulations, we set $N_{\text{V}_{\text{O}}^{..},\text{c}}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}$, $N_{\text{dop},\text{c}}$ and $N_{\text{dop},\text{b}}$ as $1\times 10^{27}$ $\text{m}^{-3}$, $5.12\times 10^{28}\text{m}^{-3}$, $1.68\times 10^{27}$ $\text{m}^{-3}$ and $1.69\times 10^{28}$ $\text{m}^{-3}$. The simulations are conducted at a temperature of 1623 K, where the non-Arrhenius behavior of the grain growth is found [28]. Periodic boundary conditions are applied to all edges of the simulation domain for the phase-fields and concentration fields. For the electrostatic potential field, the bottom-left corner is grounded, while the other edges are set with the periodic boundary conditions. Furthermore, all the parameters used in the simulation are normalized. The lattice constant $a$ = $3.9\times 10^{-10}$ m is used for the length normalization. Other normalization factors include e.g. $f_{0}=\kappa/a^{2}$=$2.3\times 10^{9}$ $\text{J}\text{m}^{-3}$ for the energy density, $t_{0}=1/(Lf_{0})=0.0043$ s for the interface mobility, and $\phi_{0}=f_{0}V_{\text{m}}/\mathcal{F}$= 0.858 V for the electrostatic potential.
We present first the microstructure evolution of the grain growth process under specific conditions $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV, $\Delta g_{\text{dop}}$=-0.4 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$ when the sintering time is $t/\text{h}=\{0.067,0.33,0.67,1.0\}$. The average area of grain at time $t=0$ is <S>, i.e., 152.1 $\text{nm}^{2}$. When $t$=0.067 h, 0.33 h and 0.67 h, the grain size increases while the area fraction of small grain decreases. The grain size distribution starts to transition from a mono-modal to a bimodal pattern. When $t$=1 h, the largest grain area is about 32 times larger than <S>. The bimodal grain size distribution is clearly visible.
Figure 6: The microstructure evolution of grain growth process at (a) 0.067 h, (b) 0.33 h, (c) 0.67 h and (d) 1 h, with defect chemistry parameters set at $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV, $\Delta g_{\text{dop}}$=-0.4 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$. These parameters are consistent with those used in Fig. 7(e). By 1h, a bimodal grain size distribution becomes apparent.
@@ 5.3.1 Influence of segregation energies and bulk defect concentration
Subsequently, we investigate the influence of defect chemistry parameters on the microstructure evolution during sintering processes. The microstructure of grains is presented in Fig. 7(a)-(g) after 1 hour of sintering at 1623 K. In Fig. 7(a), (b) and (c), we increase $\Delta g_{\text{V}_{\text{O}}^{..}}$ from -1.5 eV to -0.5 eV when $\Delta g_{\text{dop}}$=-0.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$. Bimodal grain size distribution is clearly observed in Fig. 7(a2), (b2) and (c2). As $\Delta g_{\text{V}_{\text{O}}^{..}}$ becomes more negative, the peak in the range of $40<\frac{S}{<S>}<$ 4 induced by smaller grains becomes more pronounced. As predicted in Fig. 5(a), solute drag effect is almost negligible for $\Delta g_{\text{V}_{\text{O}}^{..}}$=-0.5 eV. Nonetheless, the formation of solute cloud imprinted by vanished GBs is still detectable in Fig. 7(c3) even though the solute drag effect is very weak. Concentration of oxygen vacancy is detailed in Fig. 7(a4), (b4) and (c4). Segregation is more significant for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV. Compared to the dopant, solute cloud of oxygen vacancies are less distinct. In view of the higher diffusivity of oxygen vacancy, the contribution of oxygen vacancy to solute drag force is also weak. Solute clouds of oxygen vacancies are only observed near the junctions where grains shrink rapidly with a small radius. The electrostatic potential distributions are depicted in Fig. 7(a5), (b5) and (c5). The negative potential bands can be also observed near the GB cores. As all the considered cases show, the electrostatic potentials in small shrinking grains are significantly higher than in the surrounding growing grains, leading to a stronger electrostatic contribution to the solute drag effects. Consequently, the shrinking speed of small grains is reduced, which induces pinning effects on the surrounding growing grains. Note that the heterogeneous electrostatic potential distribution arises from the velocity difference of individual GB, in accordance to the observation in the quasi-static cases in Fig. 4, rather than any anisotropy which is absent in the current simulations.
@@
In Fig. 7(a), (d) and (e), we keep $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $c_{\text{dop},\text{b}}$ at -1.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$ and test different $\Delta g_{\text{dop}}$ values. Whereas, in Fig. 7(a), (f) and (g), $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ are constant while $c_{\text{dop},\text{b}}$ increases from $2\times 10^{25}$ $\text{m}^{-3}$ to $2\times 10^{26}$ $\text{m}^{-3}$. For all simulations, the solute drag effects from dopant are pronounced. The dopant clouds imprinted by vanished GBs are clearly observed behind GB cores. In Fig. 7(e2) and (g2), the bimodal grain size distributions are presented. However, when $\Delta g_{\text{dop}}=-0.6$ eV in Fig. 7(d2) and $c_{\text{dop},\text{b}}=2\times 10^{26}$ $\text{m}^{-3}$ in Fig. 7(f2), the unimodal grain size distribution exhibits, due to the more pronounced probability of small grain. The simulation results in Fig. 7(f) demonstrate the high grain number, where the large grains are surrounded by small grains.
@@
Figure 7: Phase-field simulation results of abnormal grain growth in Fe-doped STO after 1 h sintering under the influence of different defect chemistry parameters.
@@ 5.3.2 Grain boundary potential distribution
In Fig. 8, we present the statistics of the average grain boundary potential in the GB core, denoted as $\bar{\phi}_{\text{core}}$. The value of $\bar{\phi}_{\text{core}}$ is calculated as an average of the electrostatic potential along the GB core indicated by $\eta_{i}=0.5$ over individal GB length. Fig. 8(a) illustrates the distribution of $\bar{\phi}_{\text{core}}$ for $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig. 7(c) for the associated phase-field simulation result). When the solute drag effect is negligible, $\bar{\phi}_{\text{core}}$ is smaller than 0.05 V and approximately follows the normal distribution. In Fig. 8(b), the distribution of $\bar{\phi}_{\text{core}}$ is presented for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig. 7(a)). Here, $\bar{\phi}_{\text{core}}$ significantly increases. In addition, the distribution of $\bar{\phi}_{\text{core}}$ changes from mono-modal to bimodal, with one peak at 0.31 V and another at 0.35 V. In Fig. 8(c), the parameters $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=2\times 10^{26}$ $\text{m}^{-3}$ are chosen (refer to Fig. 7(f)). In this case, the solute drag effect is the most pronounced in this study. Not only $\bar{\phi}_{\text{core}}$ increases but also the bimodal distribution of $\bar{\phi}_{\text{core}}$ becomes more distinct.
@@
In addition, two distinct types of GBs can be observed Fig. 9. The first type, characterized by a lower moving velocity, exhibits nearly symmetric concentration profiles across the GB core with a significantly higher segregation concentration of dopant. However, within a polycrystalline system, the accumulation zone of dopant at the space charge region transitions into a depletion zone. Conversely, the second type of GB core, which moves at a higher velocity, shows lower dopant concentrations due to the break of the symmetric SCL. Similarly, the oxygen vacancy distribution in this second GB core is also asymmetric, as illustrated in Fig. 7(e). Fig. 7(f) presents the electrostatic potential distributions across these two types of GB cores, where the second GB core displays a larger grain boundary potential and a more pronounced electrostatic potential difference.

### v3

@@ 5.3 Polycrystalline simulations during sintering
After systematically benchmarks and knowledge gained from quasi-static single GB cases, we investigate different grain growth patterns of materials with only consideration of solute drag effects. with different defect segregation energies $\Delta g_{\text{dop}}$, $\Delta g_{\text{V}_{\text{O}}^{..}}$ and bulk defect concentrations such as $c_{\text{dop},\text{b}}$. The segregation of point defects in the core plays a significant role in influencing grain growth kinetics. To investigate the grain size distribution of STO polycrystalline material after sintering, we conduct 2D phase-field simulations. Note that three-dimensional phase-field simulations of the proposed model are achievable but require significantly more computational resources. In fact, to reveal the fundamental mechanisms, 2D simulations should be sufficient and allow for a comprehensive parameter study given the limited computational resources.
@@
The defect chemistry parameters used in this section are tabulated in Table 4. Then, we describe the FEM settings for the 2D phase-field simulations as follows. The simulation domain is 390 nm $\times$ 390 nm, with a uniform mesh spacing of $\Delta x=\Delta y=0.39$ nm. The GB core width is set to 0.78 nm, consistent with the value used in the 1D simulations to ensure accurate reproduction of the SCLs. The polycrystalline microstructure is initialized using a Voronoi tessellation method, where each grain is assigned a unique grain index. The initial configuration consists of 1000 grains, and the random seed for generating the polycrystalline microstructure is set to 42. To optimize memory usage during the simulations, the grain tracker algorithm is employed yang20193d . Periodic boundary conditions are applied to all edges of the simulation domain for both the phase-field and concentration fields. For the electrostatic potential field, the bottom-left corner is grounded, while the remaining edges are set with periodic boundary conditions. Each simulation is performed on 8 nodes (a total of 768 cores) with 3072 gigabytes of memory, requiring approximately 30 days to complete.
5.3.1 Skewed grain size distribution
Compared with normal grain growth pattern, the grain size distribution is skewed when the defect segregation in the GB core is considered as shown in Fig. 6. To clearly distinguish the grain size distribution, a histogram of grain area fraction is constructed. The bin width, denoted as $\xi_{\text{bin}}$, is determined using the Freedman-Diaconis rule freedman1981histogram . Subsequently, the grain size distribution is fitted as a sum of two weighted lognormal distribution functions vikrant2020electrochemicallyActa , given by $\text{PDF}(\hat{S})=\hat{k}_{1}\mathcal{N}_{1}(\hat{S},\hat{\mu}_{1},\hat{\sigma}_{1})+\hat{k}_{2}\mathcal{N}_{2}(\hat{S},\hat{\mu}_{2},\hat{\sigma}_{2})$, where $\hat{S}$ is the normalized grain area and can be expressed as $\hat{S}=S/\langle S\rangle/\xi_{\text{bin}}$. $\langle S\rangle=152.1$ $\text{nm}^{2}$ is the average area of grain at time $t=0$. $\hat{k}_{1}$ and $\hat{k}_{2}$ are the proportionality coefficients. $\hat{\mu}_{1}$ and $\hat{\mu}_{2}$ are the mean grain areas. $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ are the standard deviations. the lognormal distributions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ share the general form $\mathcal{N}=\frac{1}{\hat{S}\hat{\sigma}\sqrt{2\pi}}\exp\left[-\frac{(\ln\hat{S}-\hat{\mu})^{2}}{2\hat{\sigma}^{2}}\right]$.
The formation of skewed grain size distribution during the sintering process is demonstrated in Fig. 6. The dopant concentration is $2\times 10^{26}$ $\text{m}^{-3}$, corresponding to a 1.1% site fraction, with segregation energies of oxygen vacancy and acceptor dopant set to -1.5 eV and -0.5 eV, respectively. At a sintering time of 1 h, the grain size distribution is nearly monomodal, as shown in Fig. 6(a2), with only one peak observed in the fitted lognormal function. By 1.5 h, a transition from monomodal to skewed grain size distribution occurs, as illustrated in Fig. 6(b2), where two distinct peaks emerge. At 2 h and 3 h, the grain size distribution is skewed and do not follow log-normal distributions.
The peak positions of the two grain populations are shown in Fig. 7(a). For small grains, the peak position increases from 2.55 at 1.5 h to 3.07 at 3 h, while for large grains, it rises from 6.21 to 10.64 over the same period. As sintering progresses, the size disparity between small and large grains becomes increasingly significant. In Fig. 7(b), the time evolution of grain area is plotted for 10 small grains and 10 large grains. The unique grain numbers are labeled in Fig. 6(d1), with black numbers indicating small grains and red numbers representing large grains. The results reveal that both small and large grains grow rapidly during the initial stage of sintering. However, as sintering progresses, the grain areas of small grains stabilize, while those of large grains continue to grow. Therefore, the growth velocity difference between two distinct grain populations leads to the formation of a skewed grain size distribution, which do not follow the normal grain growth.
Moreover, the peak positions of the two grain populations in this skewed grain size distribution are influenced by various defect chemistry parameters, such as the segregation energies of dopant and oxygen vacancy, as well as the dopant concentration in the bulk phase. When these defect chemistry parameters enhance the solute drag effect, smaller grains become pinned earlier and remain stable during the sintering process, leading to a more noticeable difference between the two peak positions. In the following, we examine how the selection of defect chemistry parameters affects the grain growth processes.
Figure 6: The microstructure evolution during the grain growth process is presented at (a1) 1 h, (b1) 1.5 h, (c1) 2 h, and (d1) 3 h, with defect chemistry parameters set to $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, and $c_{\text{dop},\text{b}}=2\times 10^{26}\text{m}^{-3}$. These parameters are consistent with those used in Fig. 8(f). The area fraction distribution is fitted by two lognormal distribution functions and shown in (a2)–(d2) to illustrate the skewed grain size distribution. The blue dashed lines represent the smaller grain population, while the red dashed lines correspond to the larger grain population. The brown solid lines indicate the combined distribution of both populations. Figure (a) also appears in Fig. 8 (f) to compare with other simulations.
Figure 7: (a) Illustration of the grain size peak positions for two grain populations: small grains and large grains. For sintering times shorter than 1 h, a monomodal grain size distribution is observed, with only a single peak present. As sintering progresses, a skewed grain size distribution emerges, characterized by two distinct peaks. The difference in peak positions between small and large grains becomes increasingly pronounced with longer sintering times. (b) Time evolution of grain areas for 20 specific grains. The unique grain numbers are labeled in Fig. 6. Solid lines represent small grains, while dashed lines correspond to large grains. As sintering time increases, the grain areas of large grains grow significantly, while the grain areas of small grains remain pinned.
@@ 5.3.2 Influence of segregation energies and bulk defect concentration
Subsequently, we investigate the influence of defect chemistry parameters on the microstructure evolution during sintering processes. The microstructure of grains is presented in Fig. 8(a)-(g) after 1 hour of sintering at 1623 K. In Fig. 8(a), (b) and (c), we decrease the segregation energy of oxygen vacancy ($\lvert\Delta g_{\text{V}_{\text{O}}^{..}}\rvert$) from 1.5 eV to 0.5 eV when $\Delta g_{\text{dop}}$=-0.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$. Skewed grain size distribution is clearly observed in Fig. 8(a2). The reduction of oxygen vacancy segregation energy decreases not only the segregation concentration of oxygen vacancy but also that of the dopant in the GB core. Consequently, the solute drag effect is notably weakened. As shown in Fig. 8(b2), when $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.0\text{eV}$, two distinct grain populations are still observed. In contrast, Fig. 8(c2) shows that the grain size distribution becomes nearly monomodal when $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5\text{eV}$, indicating that the solute drag effect is negligible in this case. This observation is consistent with the 1D phase-field simulation results presented in Fig. 5. The concentration of dopant is presented in Fig. 8(a3), (b3) and (c3), the formation of solute cloud imprinted by vanished GBs is clearly observed, especially in Fig. 8(a3). Concentration of oxygen vacancy is detailed in Fig. 8(a4), (b4) and (c4). Segregation is more significant for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV. Compared to the dopant, solute cloud of oxygen vacancies are less distinct. In view of the higher diffusivity of oxygen vacancy, the contribution of oxygen vacancy to solute drag force is also weak. Solute clouds of oxygen vacancies are only observed near the junctions where grains shrink rapidly with a small radius. The electrostatic potential distributions are depicted in Fig. 8(a5), (b5) and (c5). The GB potential considerably decrease from (a5) to (c5). The negative potential bands can be also observed near the GB cores. As all the considered cases show, the electrostatic potentials in small shrinking grains are significantly higher than in the surrounding growing grains, leading to a stronger electrostatic contribution to the solute drag effects. Consequently, the shrinking speed of small grains is reduced, which induces pinning effects on the surrounding growing grains. Note that the heterogeneous electrostatic potential distribution arises from the velocity difference of individual GB, in accordance to the observation in the quasi-static cases in Fig. 4, rather than any anisotropy which is absent in the current simulations.
In Fig. 8(a), (d), and (e), we fix $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$, and examine the influence of different $\Delta g_{\text{dop}}$ values. A larger $\lvert\Delta g_{\text{dop}}\rvert$ leads to more pronounced solute drag effects. In all cases, a skewed grain size distribution is observed, even when $\Delta g_{\text{dop}}$ = -0.2 eV, as shown in Fig. 8(e). In Fig. 8(d2), although the solute drag effect is more pronounced due to higher dopant segregation energy and the grain size distribution can be fitted by two lognormal distribution functions, the PDF($\hat{S}$) function still exhibits a monomodal shape. We expect that a skewed PDF($\hat{S}$) function that do not follow the normal grain growth will emerge if the sintering time is extended, as demonstrated in Fig. 6. Additionally, the electrostatic potential distribution remains largely unchanged, as shown in Fig. 8(a5), (d5), and (e5). For the low dopant segregation energy case ($\Delta g_{\text{dop}}=-0.2$ eV), the solute drag effect arises primarily from the electrochemical potential gradient terms (see Eq. 65), which is contributed by the electrostatic potential part, rather than chemical part.
@@
In Fig. 8(a), (f) and (g), $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ are constant while $2\times 10^{25}$ $\text{m}^{-3}$, $7\times 10^{25}$ $\text{m}^{-3}$ and $2\times 10^{26}$ $\text{m}^{-3}$ are chosen for dopant concentration in bulk phase, corresponding to 0.11%, 0.42% and 1.1% site fractions. For all simulations, the solute drag effects from dopant are pronounced. The dopant clouds imprinted by vanished GBs are clearly observed behind GB cores. In Fig. 8(e2) and (g2), the skewed grain size distributions are presented. However, due to the more pronounced probability of small grain, longer sintering time is required for $c_{\text{dop},\text{b}}(\infty)$ = $2\times 10^{26}$ $\text{m}^{-3}$ to exhibit the skewed grain size distribution that do not follow the normal grain size distribution (also see Fig. 6).
The influence of different defect chemistry parameters on grain growth process are elucidated in this part. First, the segregation oxygen vacancy plays a dominant role in solute drag induced skewed grain size distribution. Reduction of the segregation energy of oxygen vacancy is an effective way to eliminate solute drag effects and the skewed grain size distribution to enhance the sintering capability of oxide ceramics. In our case, when $\lvert\Delta g_{\text{V}_{\text{O}}^{..}}\rvert$< 0.5 eV, the skewed grain growth behavior can be nearly avoided. Second, the segregation energy of dopant impacts the SCL formation and GB concentration of dopant. Increase segregation energy of dopant strongly decrease the grain growth rate. However, very low $\lvert\Delta g_{\text{dop}}\rvert$< still leads to skewed grain size distribution. Third, dopant concentration dependent solute drag effect is also verified in the present work, even very small dopant concentration (0.11%) can lead to the skewed grain size distribution that do not follow log-normal distributions.
@@
Figure 8: Phase-field simulation results of different grain growth patterns in Fe-doped STO after 1 h sintering under the influence of different defect chemistry parameters.
@@ 5.3.3 Grain boundary potential distribution
In Fig. 9, we present the statistics of the average grain boundary potential in the GB core, denoted as $\bar{\phi}_{\text{core}}$. The value of $\bar{\phi}_{\text{core}}$ is calculated as an average of the electrostatic potential along the GB core indicated by $\eta_{i}=0.5$ over individal GB length. Fig. 9(a) illustrates the distribution of $\bar{\phi}_{\text{core}}$ for $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig. 8(c) for the associated phase-field simulation result). When the solute drag effect is negligible, $\bar{\phi}_{\text{core}}$ is smaller than 0.05 V and approximately follows the normal distribution. In Fig. 9(b), the distribution of $\bar{\phi}_{\text{core}}$ is presented for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig. 8(a)). Here, $\bar{\phi}_{\text{core}}$ significantly increases. In addition, the distribution of $\bar{\phi}_{\text{core}}$ changes from mono-modal to skewed, with one peak at 0.31 V and another at 0.35 V. In Fig. 9(c), the parameters $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=2\times 10^{26}$ $\text{m}^{-3}$ are chosen (refer to Fig. 8(f)). In this case, the solute drag effect is the most pronounced in this study. Not only $\bar{\phi}_{\text{core}}$ increases but also the skewed distribution of $\bar{\phi}_{\text{core}}$ becomes more distinct.
@@
In addition, two distinct types of GBs can be observed Fig. 10. The first type, characterized by a lower moving velocity, exhibits nearly symmetric concentration profiles across the GB core with a significantly higher segregation concentration of dopant. However, within a polycrystalline system, the accumulation zone of dopant at the space charge region transitions into a depletion zone. Conversely, the second type of GB core, which moves at a higher velocity, shows lower dopant concentrations due to the break of the symmetric SCL. Similarly, the oxygen vacancy distribution in this second GB core is also asymmetric, as illustrated in Fig. 8(e). Fig. 8(f) presents the electrostatic potential distributions across these two types of GB cores, where the second GB core displays a larger grain boundary potential and a more pronounced electrostatic potential difference. These two distinct types of GBs play different roles in the grain growth process. Fast moving GBs facilitate the formation of the large grain population during sintering, while low-moving GBs are responsible for pinning the small grain population.
@@
In summary, we demonstrates the formation of a skewed grain size distribution in Fe-doped SrTiO3 during the sintering process, driven solely by the solute drag effect. Two types of GBs emerge during the growth process due to defect segregation in the GB core: fast-moving GBs and slow-moving GBs. Fast-moving GBs facilitate the growth of large grains, while slow-moving GBs act to pin small grains. The coexistence of these GBs ultimately leads to the skewed grain size distribution. Furthermore, variations in defect chemistry significantly influence the resulting grain size distribution. To improve the sintering process, reducing the segregation energy of oxygen vacancies proves to be an effective approach. Lowering the dopant segregation energy and bulk concentration also mitigate solute drag effects. However, even at low dopant segregation energy and bulk dopant concentration, the skewed grain size distribution can still occur.